# OccMamba: Semantic Occupancy Prediction with State Space Models

### v1

@@ Abstract
Training deep learning models for semantic occupancy prediction is challenging due to factors such as a large number of occupancy cells, severe occlusion, limited visual cues, complicated driving scenarios, etc. Recent methods often adopt transformer-based architectures given their strong capability in learning input-conditioned weights and long-range relationships. However, transformer-based networks are notorious for their quadratic computation complexity, seriously undermining their efficacy and deployment in semantic occupancy prediction. Inspired by the global modeling and linear computation complexity of the Mamba architecture, we present the first Mamba-based network for semantic occupancy prediction, termed OccMamba. However, directly applying the Mamba architecture to the occupancy prediction task yields unsatisfactory performance due to the inherent domain gap between the linguistic and 3D domains. To relieve this problem, we present a simple yet effective 3D-to-1D reordering operation, i.e., height-prioritized 2D Hilbert expansion. It can maximally retain the spatial structure of point clouds as well as facilitate the processing of Mamba blocks. Our OccMamba achieves state-of-the-art performance on three prevalent occupancy prediction benchmarks, including OpenOccupancy, SemanticKITTI and SemanticPOSS. Notably, on OpenOccupancy, our OccMamba outperforms the previous state-of-the-art Co-Occ by 3.1% IoU and 3.2% mIoU, respectively. Codes will be released upon publication.
@@ 1 Introduction
Figure 1: (a) Challenges in semantic occupancy prediction, (b) comparison in GPU memory consumption, and (c) performance comparison on OpenOccupancy validation set. Our OccMamba demonstrates high efficiency in handling a large number of voxel grids, outperforming all other semantic occupancy predictors, such as TPVFormer (Huang et al. 2023) and M-CONet (Wang et al. 2023b) under the settings of RGB-only, LiDAR-only and multi-modal fusion.
* The corresponding authors.
@@
Semantic occupancy prediction is becoming an indispensable component in autonomous driving, augmented reality, robotics (Roldao, De Charette, and Verroust-Blondet 2022), etc., which estimates the occupancy and categorical labels of the surrounding environment. It faces many challenges, such as the enormous number of occupancy grids, severe occlusion, limited visual clues as well as complex driving scenarios (Zhang et al. 2024b).
@@
Recent attempts, such as MonoScene (Cao and De Charette 2022) and JS3C-Net (Yan et al. 2021), have made progress in addressing these challenges, but limitations remain due to their reliance on uni-modal inputs. Multi-modal approaches, such as FusionOcc (Zhang et al. 2024a) and M-CONet (Wang et al. 2023b), offer improvements, yet they struggle to capture global information due to the inherent deficiency of CNN architectures. Despite the success of transformer-based models (Li et al. 2023b; Tong et al. 2023; Zhang, Zhu, and Du 2023), they suffer from high computational complexity, particularly when processing a large number of voxel grids.
@@
Mamba (Gu and Dao 2023), which is an important variant of state space models, emerges as a promising next-generation structure for replacing the transformer architecture. In semantic occupancy prediction, we anticipate leveraging its global modeling capabilities to better handle the complex scenarios and overcome limited visual cues, while its linear computation complexity helps manage the enormous number of occupancy grids efficiently. However, it is primarily designed for language modeling, with 1D input data, whereas the input data for semantic occupancy prediction is 3D voxels. The use of dense voxels in large scenes makes the deployment of Mamba even more challenging. When converting 3D data into a 1D format, the inherent spatial relationships between adjacent voxels in the 3D space are missing, causing neighboring voxels to become far apart in the 1D sequence. This spatial separation undermines Mamba’s ability to effectively understand local and global scene contexts, hampering the prediction performance. To address this issue, developing an effective reordering strategy is crucial to preserving spatial proximity in the transformation from 3D to 1D. Several attempts have been made to the design of such a reordering policy. For instance, Point Mamba (Liang et al. 2024) rearranges the point clouds according to the 3D Hilbert curve (Hilbert 2013) and concatenates the features of the reordered points. While these advancements have shown potential in the point cloud classification and segmentation field, the exploration of Mamba-based architectures is still in its infancy in outdoor semantic occupancy prediction tasks.
@@
In this work, we present the first Mamba-based network for semantic occupancy prediction, which we refer to as OccMamba. Owing to the global modeling and linear computation complexity of Mamba, our OccMamba can efficiently process a large number of voxel grids given limited computation resources, as shown in Fig. 1(a)-(b). To facilitate the processing of Mamba blocks as well as preserve the spatial structure of the 3D input data, we present a simple yet effective 3D-to-1D reordering policy, i.e., height-prioritized 2D Hilbert expansion. The designed policy sufficiently utilizes the categorical clues in the height information as well as the spatial prior in the XY plane. In this way, OccMamba effectively exploits LiDAR and camera cues, enabling effective fusion and processing of voxel features extracted from these sources without compression. This uncompressed voxel feature processing method, combined with Mamba’s global modeling capabilities, enhances OccMamba with the occlusion reasoning capability, particularly in complex driving scenarios. We perform extensive experiments on three semantic occupancy prediction benchmarks, i.e., OpenOccupancy (Wang et al. 2023b), SemanticKITTI (Behley et al. 2019) and SemanticPOSS (Pan et al. 2020). Our OccMamba consistently outperforms state-of-the-art algorithms in all benchmarks. It is noteworthy that on OpenOccupancy, partly shown in Fig. 1(c), our OccMamba surpasses the previous state-of-the-art Co-Occ (Pan, Wang, and Wang 2024) by 3.1% IoU and 3.2% mIoU, respectively.
The contributions of our paper are summarized as follows:
@@
  • To the best of our knowledge, we design the first Mamba-based network for outdoor semantic occupancy prediction. It possesses global modeling capability with linear computation complexity, which is essential to the processing of a large number of occupancy grids.
@@
  • To facilitate the processing of Mamba blocks as well as maximally retain the original spatial structure of point clouds, we design a simple yet effective reordering policy that projects the point clouds into 1D sequences.
@@ 2 Related work
Semantic Occupancy Prediction. The objective of semantic occupancy prediction is to estimate the occupancy and semantic labels of 3D spaces using various types of input signals (Roldao, De Charette, and Verroust-Blondet 2022). Given that LiDAR and RGB-D camera can provide accurate spatial measurement, many studies relied on such detailed geometric data like LiDAR points (Rist et al. 2021; Xia et al. 2023; Roldao, de Charette, and Verroust-Blondet 2020; Yan et al. 2021) and RGB-D images (Chen et al. 2020; Li et al. 2020a; Garbade et al. 2019; Li et al. 2020b). Meanwhile, image-based methods, increasingly popular due to camera accessibility, such as MonoScene (Cao and De Charette 2022) and TPVFormer (Huang et al. 2023), estimate occupancy of the environment using RGB images. However, inaccurate depth estimation in RGB images results in lower performance than LiDAR-based models. In response, multi-modal fusion combining both modalities has attracted significant attention. CONet (Wang et al. 2023b) and Co-Occ (Pan, Wang, and Wang 2024) demonstrate the advantages of combining modalities, enhancing precision and reliability. Furthermore, recent methods have shifted from CNNs to integrating transformers, as seen in OccFormer’s dual-path transformer (Zhang, Zhu, and Du 2023) and OccNet’s cascade refinement (Tong et al. 2023) reduce the transformer’s computational demands.
@@
Multi-Modal Fusion. Multi-modal fusion aims to make the strength of different input modalities, thus making more accurate and robust perception (Gao et al. 2020; Li et al. 2020a; Liu et al. 2023; Wang et al. 2023a). For instance, AICNet (Li et al. 2020a) integrates RGB and depth data using anisotropic convolutional networks to enhance the accuracy and completeness of semantic occupancy prediction. PointPainting (Vora et al. 2020) performs semantic segmentation on RGB images and attaches semantic probabilities to LiDAR points, thereby enriching the point cloud with rich semantic information. For semantic occupancy prediction, images and point clouds are two prevalent input signals. Recent trends in the occupancy prediction field favour the multi-modal fusion as it utilizes the strength of both signals.
@@
Figure 2: Schematic overview of our OccMamba. Given surround-view images and LiDAR point clouds, we first employ 2D encoder and 3D encoder to process them and obtain camera features and LiDAR voxel features, respectively. View transformer is utilized to project 2D camera features into 3D space, producing camera voxel features. Camera voxel features and LiDAR voxel features are fused and sent to the hierarchical Mamba module, where the proposed height-prioritized 2D Hilbert expansion reordering scheme is used to maximally utilize the spatial clues of voxels. Eventually, the Mamba features are fed to the occupancy head, producing semantic occupancy predictions.
@@
State Space Models. The transformer (Vaswani et al. 2017) has reshaped the computer vision field but suffers from the quadratic computation complexity (Lin et al. 2022). To relieve this, more efficient operators like linear attention (Wang et al. 2020), flash attention (Dao et al. 2022) have been proposed. State space model series (SSMs) such as Mamba (Gu and Dao 2023), S4 (Gu, Goel, and Ré 2021), and S4nd (Nguyen et al. 2022) are gaining prominence, with Mamba being notable for integrating selective mechanisms, which can effectively capture long-range dependencies and process large-scale data in linear time. This innovation has extended into computer vision domain through variants like VMamba (Liu et al. 2024b), which includes a cross-scan module, and Vision Mamba (Zhu et al. 2024), which utilizes a bidirectional SSM. In point cloud processing, PointMamba (Liang et al. 2024) improves the global modeling of point clouds by rearranging input patches based on 3D Hilbert curve, while Point Mamba (Liu et al. 2024a) employs an octree-based ordering for efficient spatial relationship capture. These developments demonstrate Mamba’s proficiency in processing large scale 3D data. Building on these advancements, we attempt to design a Mamba-based network tailored for semantic occupancy prediction.
@@ 3 Methodology
To efficiently and effectively process a large number of voxel grids in semantic occupancy prediction, we propose to take Mamba (Gu and Dao 2023) as the basic building block that enjoys the benefit of both global modeling and linear computation complexity. To facilitate the processing of Mamba blocks, we design a simple yet effective reordering scheme, dubbed height-prioritized 2D Hilbert expansion, along with the Hierarchical Mamba Encoder and Decoder. In the following sections, we first provide a brief review of state space models and Mamba in Sec. 3.1. Then, we present the framework overview of OccMamba in Sec. 3.2. Thereafter, we provide a detailed explanation of the reordering scheme, the hierarchical Mamba encoder and decoder in Sec. 3.3. Eventually, the training objective is presented in Sec. 3.4.
3.1 Preliminaries
$h_{k}$ $=\overline{A}h_{k-1}+\overline{B}x_{k},$ (1)
$y_{k}$ $=\overline{C}h_{k},$
where $k$ is the sequence number, $\overline{A}$, $\overline{B}$, $\overline{C}$ are matrices that represent the discretized parameters of the model, which involve the sampling step $\Delta$. The $x_{k}$, $y_{k}$, and $h_{k}$ denote the input, output, and hidden state of the system, respectively. Structure state space sequence models (S4) (Gu, Goel, and Ré 2021) introduces a new parameterization for SSMs to achieve more efficient computation.
Mamba Module. Mamba introduces an adaptation to the S4 models (Gu, Goel, and Ré 2021), where it makes the matrices $\overline{B}$ and $\overline{C}$, as well as the sampling size $\Delta$, dependent on the input. This dependency arises from incorporating the sequence length and batch size of the input, enabling dynamic adjustments of these matrices for each input token. By this way, Mamba allows $\overline{B}$ and $\overline{C}$ to dynamically influence the state transition conditioned on the input, enhancing the model’s content-awareness.
3.2 Framework Overview
Figure 2 depicts the pipeline of our OccMamba, which is built upon M-CONet (Wang et al. 2023b).
Multi-Modal Visual Encoders. Taking both point cloud and multi-view images as input, OccMamba processes each modality with respective visual encoder. Specifically, for the LiDAR branch, we first voxelize the input point clouds $\mathbf{P}$ (Zhou and Tuzel 2018) and then employ sparse-convolution-based LiDAR encoder $\mathbf{E}_{\mathcal{L}}$  (Yan, Mao, and Li 2018) to generate LiDAR voxel features $\mathbf{V}_{\mathcal{L}}\in\mathbb{R}^{W_{\mathcal{L}}\times H_{\mathcal{L}}\times D_{\mathcal{L}}\times C_{\mathcal{L}}}$. For the image branch, we feed the multi-view images $\mathbf{I}$ to ResNet-based image encoder (He et al. 2016) which utilizes FPN (Lin et al. 2017) to aggregate multi-scale features, and then utilize the 2D-to-3D view transformer (Liu et al. 2023) to produce image voxel features $\mathbf{V}_{\mathcal{C}}\in\mathbb{R}^{W_{\mathcal{C}}\times H_{\mathcal{C}}\times D_{\mathcal{C}}\times C_{\mathcal{C}}}$. Here, $W_{\mathcal{L}}$, $W_{\mathcal{C}}$, $H_{\mathcal{L}}$, $H_{\mathcal{C}}$, $D_{\mathcal{L}}$ and $D_{\mathcal{C}}$ are the spatial dimensions of the voxel features. $C_{\mathcal{L}}$ and $C_{\mathcal{C}}$ are the channel dimension of these voxel features. We ensure that $\mathbf{V}_{\mathcal{L}}$ and $\mathbf{V}_{\mathcal{C}}$ are identical in the spatial dimensions. Thereafter, we perform concatenation on these voxel features along the channel dimensions:
$\mathbf{V}=\text{concat}(\mathbf{V}_{\mathcal{L}},\mathbf{V}_{\mathcal{C}}).$ (2)
To process features extracted from the multi-modal encoders, we design the Mamba Encoder $\mathbf{E}_{\mathcal{M}}$ and Mamba Decoder $\mathbf{D}_{\mathcal{M}}$ as integral parts of our entire Mamba Module. In the process, voxel features $\mathbf{V}$ are fed into the Mamba module, which will be detailed in Sec. 3.3. The transformation can be described by
$\mathbf{V^{\prime}}=\mathbf{D}_{\mathcal{M}}(\mathbf{E}_{\mathcal{M}}(\mathbf{V})),$ (3)
where $\mathbf{V^{\prime}}$ retains the same size as $\mathbf{V}$.
Occupancy Head. We feed the processed features of Mamba blocks, i.e., $\mathbf{V^{\prime}}$, to the coarse-to-fine module $\mathbf{F}_{c\rightarrow f}$ (Wang et al. 2023b), which interpolates the dimensions to match the size of the ground truth, and then employ a MLP to predict the category of each voxel grid. The process is presented in the following equation:
$\mathbf{O}_{\text{occ}}=\text{MLP}(\mathbf{F}_{c\rightarrow f}(\mathbf{V^{\prime}})).$ (4)
MonoScene C 18.4 6.9 7.1 3.9 9.3 7.2 5.6 3.0 5.9 4.4 4.9 4.2 14.9 6.3 7.9 7.4 10.0 7.6
TPVFormer C 15.3 7.8 9.3 4.1 11.3 10.1 5.2 4.3 5.9 5.3 6.8 6.5 13.6 9.0 8.3 8.0 9.2 8.2
3DSketch C&D 25.6 10.7 12.0 5.1 10.7 12.4 6.5 4.0 5.0 6.3 8.0 7.2 21.8 14.8 13.0 11.8 12.0 21.2
AICNet C&D 23.8 10.6 11.5 4.0 11.8 12.3 5.1 3.8 6.2 6.0 8.2 7.5 24.1 13.0 12.8 11.5 11.6 20.2
LMSCNet L 27.3 11.5 12.4 4.2 12.8 12.1 6.2 4.7 6.2 6.3 8.8 7.2 24.2 12.3 16.6 14.1 13.9 22.2
JS3C-Net L 30.2 12.5 14.2 3.4 13.6 12.0 7.2 4.3 7.3 6.8 9.2 9.1 27.9 15.3 14.9 16.2 14.0 24.9
M-CONet C&L 29.5 20.1 23.3 13.3 21.2 24.3 15.3 15.9 18.0 13.3 15.3 20.7 33.2 21.0 22.5 21.5 19.6 23.2
Co-Occ C&L 30.6 21.9 26.5 16.8 22.3 27.0 10.1 20.9 20.7 14.5 16.4 21.6 36.9 23.5 25.5 23.7 20.5 23.5
OccMamba (ours) C 21.7 13.2 13.2 7.4 15.3 17.6 6.1 10.3 10.3 6.5 6.2 13.3 32.9 20.5 19.8 17.4 4.8 8.7
OccMamba (ours) L 36.4 22.7 26.8 11.3 20.8 26.1 14.6 16.3 20.3 14.0 17.5 20.3 39.5 24.9 25.9 25.3 28.3 30.6
OccMamba (ours) C&L 33.7 25.1 29.6 20.2 25.7 28.5 16.7 25.0 23.2 19.9 20.3 24.5 36.1 25.3 25.1 24.8 27.7 28.9
Table 1: Quantitative comparisons on OpenOccupancy validation set. C, D, L denote camera, depth and LiDAR, respectively. The best and second-best are in bold and underlined, respectively.
3.3 Hierarchical Mamba Module with Height-prioritized Reordering

### v2

@@ Abstract
This work was supported by the National Natural Science Foundation of China (No. 62332016) and the Key Research Program of Frontier Sciences, CAS (No. ZDBS-LY-JSC001).* The corresponding authors.
Training deep learning models for semantic occupancy prediction is challenging due to factors such as a large number of occupancy cells, severe occlusion, limited visual cues, complicated driving scenarios, etc. Recent methods often adopt transformer-based architectures given their strong capability in learning input-conditioned weights and long-range relationships. However, transformer-based networks are notorious for their quadratic computation complexity, seriously undermining their efficacy and deployment in semantic occupancy prediction. Inspired by the global modeling and linear computation complexity of the Mamba architecture, we present the first Mamba-based network for semantic occupancy prediction, termed OccMamba. Specifically, we first design the hierarchical Mamba module and local context processor to better aggregate global and local contextual information, respectively. Besides, to relieve the inherent domain gap between the linguistic and 3D domains, we present a simple yet effective 3D-to-1D reordering scheme, i.e., height-prioritized 2D Hilbert expansion. It can maximally retain the spatial structure of 3D voxels as well as facilitate the processing of Mamba blocks. Endowed with the aforementioned designs, our OccMamba is capable of directly and efficiently processing large volumes of dense scene grids, achieving state-of-the-art performance across three prevalent occupancy prediction benchmarks, including OpenOccupancy, SemanticKITTI, and SemanticPOSS. Notably, on OpenOccupancy, our OccMamba outperforms the previous state-of-the-art Co-Occ by 5.1% IoU and 4.3% mIoU, respectively. Our implementation is open-sourced and available at: https://github.com/USTCLH/OccMamba.
@@ 1 Introduction
Figure 1: (a) Challenges in semantic occupancy prediction, (b) comparison in GPU memory consumption, and (c) performance comparison on OpenOccupancy validation set. Our OccMamba demonstrates high efficiency in handling a large number of voxel grids, outperforming all other semantic occupancy predictors, such as Co-Occ [31] and M-CONet [45].
@@
Semantic occupancy prediction is becoming an indispensable component in autonomous driving, augmented reality, robotics, etc., which estimates the occupancy and categorical labels of the surrounding environment [36]. It faces many challenges, such as the enormous number of occupancy grids, severe occlusion, limited visual clues as well as complex driving scenarios [53].
@@
Recent attempts, such as MonoScene [4] and JS3C-Net [49], have made progress in addressing these challenges, but limitations remain due to their reliance on uni-modal inputs. Multi-modal approaches, such as FusionOcc [51] and M-CONet [45], offer improvements, yet they struggle to capture global information due to the inherent deficiency of CNN architectures. Despite the success of transformer-based models [20, 40, 52], they suffer from high computational complexity, particularly when processing a large number of voxel grids.
@@
Mamba [9], which is an important variant of state space models, emerges as a promising next-generation structure for replacing the transformer architecture. In semantic occupancy prediction, we anticipate leveraging its global modeling capabilities to better handle the complex scenarios and overcome limited visual cues, while its linear computation complexity helps manage the enormous number of occupancy grids efficiently. However, it is primarily designed for language modeling, with 1D input data, whereas the input data for semantic occupancy prediction is 3D voxels. The use of dense voxels in large scenes makes the deployment of Mamba even more challenging. When converting 3D data into a 1D format, the inherent spatial relationships between adjacent voxels in the 3D space are missing, causing neighboring voxels to become far apart in the 1D sequence. This spatial separation undermines Mamba’s ability to effectively understand local and global scene contexts, hampering the prediction performance. To address this issue, developing an effective reordering strategy is crucial to preserving spatial proximity in the transformation from 3D to 1D. Several attempts have been made to achieve this, for instance, Point Mamba [21] rearranges the point clouds according to the 3D Hilbert curve [12] and concatenates the features of the reordered points. While these advancements have shown potential in the point cloud classification and segmentation field, the exploration of Mamba-based architectures is still in its infancy in outdoor semantic occupancy prediction tasks.
@@
In this work, we present the first Mamba-based network for semantic occupancy prediction, which we refer to as OccMamba. Owing to the global modeling and linear computation complexity of Mamba, our OccMamba can efficiently process a large number of voxel grids given limited computation resources, as shown in Fig. 1(a)-(b). To effectively utilize the global and local information hidden in the input voxel grids, we design the hierarchical Mamba module and local context processor. To facilitate the processing of Mamba blocks as well as preserve the spatial structure of the 3D input data, we present a simple yet effective 3D-to-1D reordering policy, i.e., height-prioritized 2D Hilbert expansion. The designed policy sufficiently utilizes the categorical clues in the height information as well as the spatial prior in the XY plane. In this way, OccMamba effectively exploits LiDAR and camera cues, enabling effective fusion and processing of voxel features extracted from these sources without compression. This uncompressed voxel feature processing method, combined with Mamba’s global modeling capabilities, enhances OccMamba with the occlusion reasoning capability, particularly in complex driving scenarios. We perform extensive experiments on three semantic occupancy prediction benchmarks, i.e., OpenOccupancy [45], SemanticKITTI [1] and SemanticPOSS [33]. Our OccMamba consistently outperforms state-of-the-art algorithms in all benchmarks. It is noteworthy that on OpenOccupancy, partly shown in Fig. 1(c), our OccMamba surpasses the previous state-of-the-art Co-Occ [31] by 5.1% IoU and 4.3% mIoU, respectively.
The contributions are summarized as follows:
@@
  • To our knowledge, we design the first Mamba-based network for outdoor semantic occupancy prediction. It possesses global modeling capability with linear computation complexity, which is crucial to the processing of a large number of voxel grids. We design the hierarchical Mamba module and local context processor to better aggregate global and local contextual information, respectively.
@@
  • To facilitate the processing of Mamba blocks as well as maximally retain the original spatial structure of 3d voxels, we design a simple yet effective reordering policy that projects the point clouds into 1D sequences.
@@ 2 Related work
Semantic occupancy prediction. The objective of semantic occupancy prediction is to estimate the occupancy and semantic labels of 3D spaces using various types of input signals [36, 13]. Given that LiDAR and RGB-D camera can provide accurate spatial measurement, many studies relied on such detailed geometric data like LiDAR points [34, 48, 35, 49] and RGB-D images [5, 15, 8, 16]. Meanwhile, image-based methods, increasingly popular due to camera accessibility, such as MonoScene [4] and TPVFormer [14], estimate occupancy of the environment using RGB images. However, inaccurate depth estimation in RGB images results in lower performance than LiDAR-based models. In response, multi-modal fusion combining both modalities has attracted significant attention. CONet [45] and Co-Occ [31] demonstrate the advantages of combining modalities, enhancing precision and reliability. Furthermore, recent methods have shifted from CNNs to integrating transformers, as seen in OccFormer’s dual-path transformer [52] and OccNet’s cascade refinement [40] reduce the transformer’s computational demands.
@@
Multi-modal fusion. Multi-modal fusion aims to make the strength of different input modalities, thus making more accurate and robust perception [7, 15, 27, 43, 18, 25, 17]. For instance, AICNet [15] integrates RGB and depth data using anisotropic convolutional networks to enhance the accuracy and completeness of semantic occupancy prediction. PointPainting [42] performs semantic segmentation on RGB images and attaches semantic probabilities to LiDAR points, thereby enriching the point cloud with rich semantic information. For semantic occupancy prediction, images and point clouds are two prevalent input signals. Recent trends in the occupancy prediction field favour the multi-modal fusion as it utilizes the strength of both signals.
@@
Figure 2: Schematic overview of our OccMamba. Given surround-view images and LiDAR point clouds, we first employ the 2D encoder and 3D encoder to process them, obtaining camera features and LiDAR voxel features, respectively. View transformer is utilized to project camera features to camera voxel features. The camera and LiDAR voxel features are then fused and sent to the hierarchical Mamba module, where the proposed height-prioritized 2D Hilbert expansion reordering is used to maximally utilize the spatial clues of voxels. Besides, local context processor is designed to divide the Mamba features into multiple windows along the XY plane, further enhancing the local semantic information. Eventually, the Mamba features are fed to the occupancy head, producing semantic occupancy predictions.
@@
State space models. The transformer [41] has reshaped the computer vision field but suffers from the quadratic computation complexity [22]. To relieve this, more efficient operators like linear attention [44], flash attention [6] have been proposed. State Space Models (SSMs) such as Mamba [9], S4 [10], and S4nd [30] are gaining prominence, with Mamba being notable for integrating selective mechanisms, which can effectively capture long-range dependencies and process large-scale data in linear time. This innovation has extended into computer vision domain through variants like VMamba [26], which includes a cross-scan module, and Vision Mamba [55], which utilizes a bidirectional SSM. In point cloud processing, PointMamba [21] improves the global modeling of point clouds by rearranging input patches based on 3D Hilbert curve, while Point Mamba [24] employs an octree-based ordering for efficient spatial relationship capture. They demonstrate Mamba’s proficiency in processing large scale 3D data. Building on these advancements, we attempt to design a Mamba-based network tailored for semantic occupancy prediction.
@@ 3 Methodology
To efficiently and effectively process a large number of voxel grids in semantic occupancy prediction, we propose to take Mamba [9] as the basic building block that enjoys the benefit of both global modeling and linear computation complexity. To facilitate the processing of Mamba blocks, we design a simple yet effective reordering scheme, dubbed height-prioritized 2D Hilbert expansion, along with OccMamba Encoder. In the following sections, we first provide a brief review of state space models and Mamba in Sec. 3.1. Then, we present the framework overview of OccMamba in Sec. 3.2. Thereafter, we provide a detailed explanation of the reordering scheme, the hierarchical Mamba module and local context processor in Sec. 3.3. Eventually, the training objective is presented in Sec. 3.4.
3.1 Preliminaries
$h_{k}$ $=\overline{A}h_{k-1}+\overline{B}x_{k},$ (1)
$y_{k}$ $=\overline{C}h_{k},$ (2)
where $k$ is the sequence number, $\overline{A}$, $\overline{B}$, $\overline{C}$ are matrices that represent the discretized parameters of the model, which involve the sampling step $\Delta$. The $x_{k}$, $y_{k}$, and $h_{k}$ denote the input, output, and hidden state of the system, respectively. As an improvement, Structured State Space Sequence Models (S4) [10] optimize traditional SSMs by introducing structured matrices, which allow the system dynamics, involving matrices $\overline{A}$, $\overline{B}$, and $\overline{C}$, to be parameterized in a way that significantly improves computational efficiency and scalability for large sequences. By leveraging these structured matrices, S4 achieves reduced computational complexity without sacrificing accuracy.
Mamba Module. Mamba introduces an adaptation to the S4 models [10], where it makes the matrices $\overline{B}$ and $\overline{C}$, as well as the sampling size $\Delta$, dependent on the input. This dependency arises from incorporating the sequence length and batch size of the input, enabling dynamic adjustments of these matrices for each input token. By this way, Mamba allows $\overline{B}$ and $\overline{C}$ to dynamically influence the state transition conditioned on the input, enhancing the model’s content-awareness. Additionally, Mamba incorporates optimizations for the scan operation and a hardware-aware algorithm, enabling efficient parallel computation.
3.2 Framework overview
Figure 2 depicts the pipeline of our OccMamba, which is built upon M-CONet [45].
Multi-modal visual encoders. Taking both point cloud and multi-view images as input, OccMamba processes each modality with respective visual encoder. Specifically, for the LiDAR branch, we first voxelize the input point clouds $\mathbf{P}$ [54] and then employ sparse-convolution-based LiDAR encoder $\mathbf{E}_{\mathcal{L}}$  [50] to generate LiDAR voxel features $\mathbf{V}_{\mathcal{L}}\in\mathbb{R}^{B\times W_{\mathcal{L}}\times H_{\mathcal{L}}\times D_{\mathcal{L}}\times C_{\mathcal{L}}}$. For the image branch, we feed the multi-view images $\mathbf{I}$ to ResNet-based image encoder [11] which utilizes FPN [23] to aggregate multi-scale features, and then utilize the 2D-to-3D view transformer [27] to produce image voxel features $\mathbf{V}_{\mathcal{C}}\in\mathbb{R}^{B\times W_{\mathcal{C}}\times H_{\mathcal{C}}\times D_{\mathcal{C}}\times C_{\mathcal{C}}}$. Here, $B$ is the batch size, $W_{\mathcal{L}}$, $W_{\mathcal{C}}$, $H_{\mathcal{L}}$, $H_{\mathcal{C}}$, $D_{\mathcal{L}}$ and $D_{\mathcal{C}}$ are the spatial dimensions of the voxel features. $C_{\mathcal{L}}$ and $C_{\mathcal{C}}$ are the channel dimension of these voxel features. We ensure that $\mathbf{V}_{\mathcal{L}}$ and $\mathbf{V}_{\mathcal{C}}$ are identical in the spatial dimensions. Thereafter, these voxel features are concatenated along the channel dimensions:
$\mathbf{V}_{\mathcal{F}}=\text{concat}(\mathbf{V}_{\mathcal{L}},\mathbf{V}_{\mathcal{C}}).$ (3)
OccMamba encoder. To process features extracted from the multi-modal encoders, we designed the hierarchical Mamba module and local context processor. The former includes the Mamba encoder $\mathbf{E}_{\mathcal{M}}$ and Mamba decoder $\mathbf{D}_{\mathcal{M}}$, while the latter is denoted as $\mathbf{P}_{\mathcal{L}}$. In the process, fusion voxel features $\mathbf{V}_{\mathcal{F}}$ are fed into them, which will be detailed in Sec. 3.3. The transformation can be described by
$\mathbf{V}_{\mathcal{M}}$ $=\mathbf{D}_{\mathcal{M}}(\mathbf{E}_{\mathcal{M}}(\mathbf{V}_{\mathcal{F}})),$ (4)
$\mathbf{V}_{\mathcal{P}}$ $=\mathbf{P}_{\mathcal{L}}(\mathbf{V}_{\mathcal{M}}),$ (5)
where $\mathbf{V}_{\mathcal{M}}$ and $\mathbf{V}_{\mathcal{P}}$ retains the same size as $\mathbf{V}_{\mathcal{F}}$.
Occupancy head. We feed the processed features of OccMamba Encoder, i.e., $\mathbf{V}_{\mathcal{P}}$, to the coarse-to-fine module $\mathbf{F}_{c\rightarrow f}$ [45], which interpolates the dimensions to match the size of the ground truth, and then employ a MLP to predict the category of each voxel grid. The process is presented in the following equation:
$\mathbf{O}_{\text{occ}}=\text{MLP}(\mathbf{F}_{c\rightarrow f}(\mathbf{V}_{\mathcal{P}})).$ (6)
@@ 3.2 Framework overview
3.3 OccMamba encoder with height-prioritized reordering